∫ Sec^6 t dt from (-π/4) to (π/4)


1
Expert's answer
2021-04-20T16:25:30-0400

We define 

"ǀ_n=∫_{-π/4}^{π/4}sec^6 t dt =[\frac {tan tsec^{n-2} t }{n-1}] ^{π/4}_{-π/4} + \frac {n-2}{n-11}ǀ_{n-2},n≥2"

As the reduction formula for the integral

"\u061e \u01c0_6=\u222bsec^6 t dt"

"=[\frac {tan t sec ^4 t}{5}]_{-π/4} ^{π/4} + \frac 45 ǀ_4…(i)"


"ǀ_4=[\frac {tan t sec ^2 t}{3}]^{\pi/4} _{-\pi/4}+ \frac 23 ǀ_2…(ii)"

"\u01c0_2= [ tan t]^{\u03c0\/4} _{-\u03c0)\/4}\u2026(iii)"

Replacing (iii) in (ii) and replacing (ii) in (i) gives

"ǀ_6=[\frac 15 tan t sec^4 t+\frac 4{15} tan t sec ^2 t+ \frac 8{15} tan t]^{π/4} _{-π)/4}"

"∫_{-π)/4}^{π/4}sec^6 t dt= [\frac15 (1) \sqrt2) ^4 + \frac 4{15} (1) \sqrt2)^2+ \frac 8{15} (-1)] - [\frac 15 (-1) sqrt2^4+ \frac 4{15} (-1)\sqrt 2^2 )+\frac 8{15 }(-1)]

=\frac {56}{15}=3.73"